### v1

# TTQA-RS- A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization

Abstract
Question answering (QA) over tables and text has gained much popularity over the years. Multi-hop table-text QA requires multiple hops between the table and text, making it a challenging QA task. Although several works have attempted to solve the table-text QA task, most involve training the models and requiring labeled data. In this paper, we have proposed a model - “TTQA-RS: A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization” [^1] Link to code in Github: https://github.com/jayetri/TTQA-RS. Our model uses augmented knowledge including table-text summary with decomposed sub-question with answer for a reasoning-based table-text QA. Using open-source language models our model outperformed all existing prompting methods for table-text QA tasks on existing table-text QA datasets like HybridQA and OTT-QA’s development set. Our results are comparable with the training-based state-of-the-art models, demonstrating the potential of prompt-based approaches using open-source LLMs. Additionally, by using GPT-4 with LLaMA3-70B, our model achieved state-of-the-art performance for prompting-based methods on multi-hop table-text QA.
TTQA-RS- A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization
Jayetri Bardhan, Bushi Xiao, Daisy Zhe Wang Department of Computer and Information Science and Engineering University of Florida Gainesville, Florida, USA {jayetri.bardhan, xiaobushi, daisyw}@ufl.edu
1 Introduction
Question Answering over tables involves extracting the table cell containing the answer to the question. The most popular approach of table QA is to generate SQL queries using the question, i.e. the table-QA task is converted into a text-to-SQL task (Pasupat and Liang, 2015; Yu et al., 2018; Zhong et al., 2017). The SQL queries are then used to retrieve the answer from the tables. Some other recent approaches use an intermediate pre-training method on the flattened tables for QA (Herzig et al., 2020; Yin et al., 2020). QA over table and text is more challenging. Datasets like HybridQA (Chen et al., 2020b) and OTT-QA (Chen et al., 2020a) are examples of multi-hop table-text QA datasets where the answer to the question can exist in the table or the text. These two datasets make use of Wikitables along with text from Wikipedia to answer the questions. The tables in the HybridQA dataset contain hyperlinks linking the table cells to Wikipedia’s text, making QA tasks more challenging. Additionally, HybridQA and OTT-QA are both multi-hop table-text datasets, which means that one or more hops between the table and text are required to derive the answer.
Figure 1: Comparison between Standard prompting, Chain of Thought prompting, and the TTQA-RS model.
Over the years, several works have attempted to solve this task. But the majority of these works have used supervised-training, requiring a large amount of labeled data (Chen et al., 2020b; Sun et al., 2021; Wang et al., 2022; Eisenschlos et al., 2021; Feng et al., 2022; Kumar et al., 2023; Chen et al., 2020a; Li et al., 2021). In this paper, we have proposed a prompting-based approach while using open-source large language models (LLMs) for multi-hop table-text QA.
With the emergence of new generative-based LLM models, prompt-based methods using in-context learning have started being explored (Chen, 2023). Training models from scratch or even fine-tuning the models requires a large amount of labeled data. In-context learning is a cheaper alternative approach that does not need any fine-tuning but instead uses pre-trained language models (LLMs) to solve new tasks using a few examples as part of the prompt. The release of the new openAI models such as GPT 4 has opened new avenues of research in natural language processing and has encouraged further research in prompt learning. (Wei et al., 2022) has shown that reasoning with chain of thought (CoT) can significantly improve the ability of large language models to perform complex reasoning in tasks including QA. But small LLMs, i.e. models with less than 100B parameters using CoT prompting tend to hallucinate and produce incorrect results, urging research communities to use bigger LLMs which are expensive and also not open-source.
In this paper, we introduced a framework - TTQA-RS, a reasoning-based prompting approach for table-text QA that despite CoT’s shortcomings on small-parameter models, we were able to reduce the hallucinations on open-source small models (i.e. we obtained a 6% increase in exact match score compared to the baseline CoT model for the HybridQA’s test set). Furthermore, our proposed model was able to beat the state-of-the-art model - S3HQA’s CoT prompting with GPT 3.5 results (Lei et al., 2023) on HybridQA dataset. By beating their model’s performance, we have shown the potential for smaller LLMs in multi-hop table-text QA.
For our experiments, we have used HybridQA dataset and OTT-QA’s development set. OTT-QA is an extension of the HybridQA dataset. Similar to the HybridQA dataset, the OTT-QA dataset is also constructed using questions based on Wikipedia tables and text. But unlike the HybridQA dataset, the test set of the OTT-QA dataset does not have hyperlinks in the table cells that can be linked to the Wikipedia text. Hence, the OTT-QA’s test set is more challenging. Existing models including ours use a retriever-reader framework for table-text QA. In this paper, we narrowed our focus to the reader of the table-text QA task. Our goal is to develop a prompting strategy for the table-text QA reader that can work even with smaller LLMs. The task of linking tables and text passages for open-domain QA is out of scope of this paper. The development set of OTT-QA, similar to the HybridQA dataset, already has hyperlinks in the table cells linking to the wiki text. In the future, we plan to extend our approach to linking the table and text for cases when hyperlinks are absent in the table cells.
The TTQA-RS model breaks down the table-text QA problem into multiple steps. In the HybridQA and the OTT-QA dataset, the questions require multiple steps of reasoning over table and text to answer. The TTQA-RS model generates the sub-questions that can help in answering the complex questions. It also generates the summary of the table and text, which is in turn used for the table-text QA of the original questions. Breaking down the complex multi-hop QA problem into simple, smaller steps can help boost the model’s overall performance. Furthermore, LLMs struggle with multi-level reasoning in a single step. So, breaking down the multi-hop QA problem along with providing an augmented information including the table-text summary can improve the performance of multi-hop table-text QA tasks using small open-source LLMs. In Figure 1, we show an example of a question from multi-hop QA that uses standard prompting, CoT, and the TTQA-RS approach for multi-hop QA.
2 Related Works
Multi-hop table-text QA can be a complex task as it requires multiple hops between the table and text to answer the questions. S3HQA (Lei et al., 2023) and MFORT-QA (Guan et al., 2024) are the only two existing models as per our knowledge that use in-context learning for multi-hop table-text QA. The S3HQA model has demonstrated table-text QA task using the Hybrid-QA dataset, whereas MFORT-QA has used the OTT-QA dataset. The S3HQA model uses a three-step method - a retriever with refinement training, a hybrid selector, and a generation-based reasoner with GPT 3.5 for the hybrid table-text QA task. MFORT-QA uses the Chain-of-thought (CoT) method to break down complex questions into smaller sub-questions, and uses Retrieval Augmented Generation to extract more context. Similar to the MFORT-QA model, we also break down complex questions into smaller sub-questions. With the complexity of the multi-hop QA task broken down into smaller questions, LLMs are in turn working on a smaller problem and perform better as single-step reasoners. Our model - TTQA-RS, additionally generates a summary using the retrieved table rows and passages. Then, for table-text question answering (QA), it uses the generated summary, the predicted entity type of the answer, and the generated sub-questions along with the answer.
Figure 2: An overview of TTQA-RS framework. The dashed lines represent the reader for the table-text QA model.
3 Our Model
3.1 System Overview
The TTQA-RS model uses a retriever-reader model. Our reader breaks down the table-text QA problem into five steps - (1) Summary generation using retrieved tables rows and passages, (2) Question decomposition, (3) Entity type prediction of the expected answer, (4) Table-text QA of independent sub-question, and (5) Table-text QA of the original question. Figure 2 shows an overview of the TTRS-QA framework. The following subsections describe the TTQA-RS framework’s retriever and reader in detail.
3.2 Retriever
The function of the retriever is to extract relevant rows and passages from the text linked to the table cells using hyperlinks. For the HybridQA dataset, we have used S3HQA model’s (Lei et al., 2023) table retriever to extract the relevant row(s) from the table, and HYBRIDER’s (Chen et al., 2020b) text-retriever to extract the relevant information from the linked passages. S3HQA’s row retriever uses refinement training to train the retriever model. The tables contain hyperlinks to Wikipedia text. So, the passages linked to the retrieved rows are collected to form a pool. The passage retriever contains an ensemble retriever of TF-IDF retriever with longest-substring retriever and selects passages with cosine distance less than a certain threshold.
For experiments on OTT-QA’s development set, we don’t use any table retriever, i.e. we only use HYBRIDER’s text retriever. The text linked to the table rows is extracted and then filtered using HYBRIDER’s text retriever.
3.3 Reader
3.3.1 Table-text Summarization
This is the first step of the reader model. The retrieved table rows are flattened with a delimiter separating the rows and columns of the table. The retrieved rows and passages are used to generate summaries of the table and text. We used zero-shot learning with LLaMA 3-70B model to generate the summaries. In Appendix B we have shown an example of a table-text summarization prompt.
Figure 3: Example of our approach using TTQA-RS model
3.3.2 Question decomposition
In the next step, we break down the questions and identify the sub-questions, such that the answer of one sub-question can aid in answering the original complex question. From here onwards, we will refer to the sub-question that can be answered first as the “independent sub-question”. Let’s take the example in Figure 3. The complex question - "What was the release date of the game which Andrew Voss provided commentary on ?" can be broken down into sub-questions. The independent sub-question for this question is - "Which game has Andrew Voss provided commentary on?". The answer to this sub-question is "Rugby League 3". This can be used to simplify the original complex question to the following - " What was the release date of Rugby League 3?". Thus, including the information about the independent sub-question and the sub-answer helps to reduce the complexity of the multi-hop task. Identifying the independent sub-question and breaking down the complex multi-hop QA problem helps to reduce the complexity of the problem, and in turn, boosts the accuracy of the model. We use in-context learning with LLaMA3-70B model to generate the independent sub-questions for the given complex queries.
3.3.3 Entity type prediction of the expected answer
We identify the entity type of the expected answer for both the independent sub-question and also for the original question. For the following question - "What was the release date of the game which Andrew Voss provided commentary on?", the entity type of the expected answer is "date". Knowing that the expected answer is of type - "date", makes the LLM’s task of generating the answer considerably easier. We have used Spacy, an open-source Python library to obtain the entity type.
3.3.4 Table-text QA of independent sub-questions
In this step, we use few-shot learning with CoT to generate the answers of the independent sub-questions. The input prompt contains the retrieved table rows, retrieved passages, the table-text summary, and also the predicted entity type of the expected answer. This is used to generate the answer for the independent sub-question.
3.3.5 Table-text QA of the original questions
This is the final step of the table-text QA framework. To generate the answers of the original questions, we use CoT-based in-context learning similar to the previous step. But in addition to the prompt containing the retrieved rows, retrieved passages, table-text summary, and the expected entity type of the predicted answer of the original question, it also includes the independent sub-question with its generated sub-answer obtained in the previous step. Figure 3 shows an example of our reader’s approach. For simplicity, we have excluded mentioning about the few-shot examples in Figure 3.
4 Experimental Setup
4.1 Datasets
HybridQA HybridQA (Chen et al., 2020b) is a large QA dataset that requires multi-hop reasoning over tables and text for QA. The questions in the HybridQA dataset are based on Wikipedia tables and corpora that are linked to the Wikipedia tables through hyperlinks.
OTT-QA (Chen et al., 2020a) is an open-domain multi-hop table-text QA dataset. For our experiments, we only use the development set of the OTT-QA dataset which contains the hyperlinks linking the table and the text (unlike OTTQA’s test set).
4.2 Implementation details
The implementation details are shown in Appendix A.
4.3 Baseline Models
Standard prompting - For the baseline standard prompting model, we used the same retriever as in TTQA-RS model, (i.e. HYBRIDER’s (Chen et al., 2020b) passage retriever with S3HQA’s (Lei et al., 2023) table retriever for the HybridQA dataset). For experiments on OTT-QA’s dev set, we don’t use any table-retriever, i.e. we only use the HYBRIDER’s passage retriever. For the reader, we performed in-context learning with standard prompting (Brown et al., 2020) for the QA task.
Chain of Thought Prompting (CoT) - Similar to the standard prompting baseline model, the CoT baseline model uses the same retriever as the TTQA-RS model. The reader uses in-context learning with CoT prompting Wei et al. (2022).
5 Results and Discussion
5.1 Main Results
Table 1: Performance of our model-TTQA-RS and other related works on the HybridQA dataset
Type Model Dev Test
EM / F1 EM / F1
Train HYBRIDER (Chen et al., 2020b) 44.0 / 50.7 43.8 / 50.6
Train DocHopper (Sun et al., 2021) 47.7 / 55.0 46.3 / 53.3
Train MuGER2 (Wang et al., 2022) 571.1 / 67.3 56.3 / 66.2
Train POINTR + MATE (Eisenschlos et al., 2021) 63.4 / 71.0 62.8 / 70.2
Train DEHG (Feng et al., 2022) 65.2 / 76.3 63.9 / 75.5
Train MITQA (Kumar et al., 2023) 65.5 / 72.7 64.3 / 71.9
Train MAFiD (Lee et al., 2023) 66.2 / 74.1 65.4 / 73.6
Train S3HQA (supervised learning) (Lei et al., 2023) 68.4 / 75.3 67.9 / 75.5
2-shot S3HQA GPT 3.5 direct (Lei et al., 2023) 57.1 / 68.8 -
2-shot S3HQA GPT 3.5 CoT (Lei et al., 2023) 60.3 / 72.1 -
2-shot Baseline Standard prompting LLaMA 3-70B 48.97 / 60.18 52.88 / 61.42
2-shot Baseline CoT LLaMA 3-70B 54.22 / 64.98 55.71 / 62.24
2-shot TTQA-RS LLaMA 3 - 70B 63.12 / 73.61 61.97 / 67.56
2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 65.49 / 76.43 63.69 / 71.83
Human - 88.2 / 93.5
(a) EM score
(b) F1 score
Figure 4: Performance of HybridQA test set on different LLaMA models
(a) EM score
(b) F1 score
Figure 5: Performance of OTT-QA dev set on different LLaMA models
Table 2: Performance of our model - TTQA-RS and other related works on OTT-QA development set.
Type Model Dev
EM F1
Train HYBRIDER (Top-1) (Chen et al., 2020b) (Chen et al., 2020b) 8.9 11.3
Train HYBRIDER (best Top-K) 10.3 13.0
Train Iterative-Retrieval + Single-Block Reader (Chen et al., 2020a) 7.9 11.1
Train Fusion-Retrieval + Single-Block Reader (Chen et al., 2020a) 13.8 17.2
Train Iterative-Retrieval + Cross-Block Reader (Chen et al., 2020a) 14.4 18.5
Train Fusion-Retrieval + Cross-Block Reader (Chen et al., 2020a) 28.1 32.5
Train CARP (Zhong et al., 2022) 33.2 38.6
Train MITQA (Kumar et al., 2023) 40.0 45.1
2-shot Baseline Standard prompting LLaMA3-70B 53.28 59.65
2-shot Baseline CoT LLaMA3-70B 55.74 63.50
2-shot TTQA-RS LLaMA 3 - 70B 63.15 70.84
2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 67.27 79.55
Table 3: Ablation studies of TTQA-RS on HybridQA and OTT-QA dataset using LLaMA 3-70B
Model HybridQA OTT-QA
Dev Test Dev
Baseline Standard prompting 48.97 / 60.18 52.88 / 61.24 53.28 / 59.65
Baseline CoT 54.22 / 64.98 55.71 / 62.24 55.74/63.50
CoT + entity-type prediction of expected answer 59.68 / 67.47 57.75 / 64.65 57.53 / 66.24
Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer with COT (no summarization) 61.14 / 70.45 58.72 / 66.40 61.23 / 69.14
Our model - Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer + summarization with CoT 63.12 / 73.61 61.97 / 67.56 63.15 / 70.84
Table 4: Human evaluation of generated summaries for a sampled test set of HybridQA
Human Evaluation Metrics Performance
Correctness 0.94
Inclusivity 0.98
Completeness 0.71
In this section, we discuss all our major findings. Table 1 displays the performance of our model with other existing models on the HybridQA dataset. We used exact match (EM) and F1 score to evaluate the performance the table-text QA models. From Table 1, we can observe that most existing models train their models for table-text QA. S3HQA (Lei et al., 2023) is the only model among the existing works that uses in-context learning for the HybridQA dataset. Please note that "S3HQA GPT-3.5 direct" refers to S3HQA model with standard prompting using GPT-3.5. Our TTQA-RS model with LLaMA 3-70B on the HybridQA’s development set was able to beat S3HQA’s CoT with GPT 3.5 by 3% exact match. Our 2-shot model with LLaMA-4 also beats the baseline standard and CoT prompting models by a huge margin (i.e. by 9% exact match when compared with standard prompting and by 6% exact match when compared with CoT in the test set). Furthermore, in Figure 4 and in Figure 5, we have shown the performance of our TTQA-RS model on different parameter models of LLaMA 2 and LLaMA 3 models on HybridQA test set and the OTT-QA development set respectively. For all the different parameter models of LLaMA-2 and LLaMA-3, our framework performed better than the baseline prompting models (i.e. standard prompting and CoT prompting). Our experiments show that our breakdown prompting approach with summarization and reasoning can improve the performance of all open-source models for table-text QA tasks. To show that our TTQA-RS approach can improve table-text QA on also GPT model, we have experimented with GPT-4 in the last stage of our model, i.e. table-text QA on the original question. For the remaining stages of the reader, we have used LLaMA 3- 70b. By adding GPT-4 in the last step, we were able to show the best performance with an exact match of 65.49 and F1 score of 76.43 in the development set, and an exact match of 63.69 and F1 score of 71.83 on the test set. With 2-shot learning using TTQA-RS LLaMA 3 70B + GPT-4 model, we were able to reach a model performance very close to the best existing training-based model (S3HQA with supervised learning) on the HybridQA dataset. For cost limitations, we have limited experiments with GPT-4 to only the last stage of our model.
Table 2 shows the performance of our model - TTQA-RS on the OTT-QA development set. To the best of our knowledge, MFORT-QA (Guan et al., 2024) is the only model that has used in-context learning for the OTT-QA dataset, but since they have not reported their performance on the development set, we therefore compare our model’s performance with other existing works that trained the models. Our TTQA-RS model with LLaMA 3-70B + GPT-4 model achieved the best performance (exact match of 67.27 and F1 score of 79.55) on the development set and has achieved new state-of-the-art performance of the OTT-QA’s development set.
With the evaluation of our model - TTQA-RS on the HybridQA and OTT-QA development set, we have shown the potential of prompting approaches with small language models (like LLaMA) and also using GPT-4.
5.2 Analysis and Ablation Studies
This section describes all the analysis and ablation studies performed on our model. Table 3 shows the ablation studies of our model using HybridQA dataset and OTT-QA’s development set. We can observe that baseline CoT model outperforms the baseline standard prompting model. This shows the importance of reasoning in the multi-hop table-text QA task. Then, we test the model by adding the entity type prediction of the expected answer in the CoT prompt. We notice a significant increase in the performance of the model for all the datasets. In the 4th row of Table 3 we have added all the components of our final model except the table-text summary and we can see a further increase in performance in the HybridQA and OTT-QA datasets. Finally, we show the performance of our model TTQA-RS (i.e. last row) by including the generated table-text summary, and we can observe that our model performs the best with all the steps included (including summarization). Adding the table-text summary in the QA input prompt, helps the LLM model to recognize relevant information related to the table or text passages that might have otherwise gone unnoticed. This shows the importance of every component of our model and the need for our break-down prompting approach for table-text QA. In Appendix  C, we evaluated the impact of the number of shots on the model’s performance.
5.3 Human Evaluation Results
We have manually evaluated the first 100 samples of the table-text summaries generated by LLaMA3-70B, and also the independent sub-questions generated using LLM prompting. We obtained an accuracy of 91% for question decomposition.
Table  4 tabulated the human evaluation results of the generated summaries for the sampled HybridQA test set. For evaluating the generated summaries using retrieved table rows and passages, we have used three evaluation metrics - correctness, inclusivity, and completeness. For correctness, we checked if the summary generated is overall correct and if the model generates any hallucination. For inclusivity, we checked if the generated summaries included information about both the retrieved rows and passages. Completeness was used to check if the generated summaries had complete sentences. We have included all our human evaluation results in the Supplementary section.
6 Conclusion
This paper proposes a prompting strategy of multi-hop table-text QA by generating table-text summaries and answers of sub-questions. We show that including summaries of retrieved table rows and passages in the prompt with our breakdown approach can substantially increase the performance of CoT prompting in table-text QA. The proposed method achieves new state-of-the-art performance among the prompting approaches for multi-hop table-text QA tasks using both open-source (i.e. LLaMA3-70B) and GPT-4 models. Our experiments specifically focussed on improving prompting strategies in the table-text QA readers. In the future, we plan to extend our work on the table-text QA retrievers which can further improve the QA performance.
Limitations
Our work has several limitations. Firstly, we are breaking down our problem into individual steps. Even though breaking down the problem into sub-problems helps to reduce hallucination while reasoning with open-source LLMs, it also causes error propagation. Errors made in the initial steps can result in wrong answers. Furthermore, in the experiment of using GPT-4 on our model, we have limited its usage only to the last step of the reader model as GPT-4 is expensive. Using GPT-4 in the remaining steps of the reader could have further improved the performance of our model.
Secondly, the performance of our prompting-based approach, even though is on par with the fine-tuned state-of-the-art models (or has outperformed the training-based state-of-the-model for OTT-QA development set), it’s performance is still not close to the human performance. Also, we are using an existing retriever and the focus of this paper has only been to improve the reader’s performance for multi-hop table-text QA. There is still potential to improve the overall performance of the model by using a better table and text retriever for this problem. Also, currently, we have only experimented with multi-hop table-text datasets in which the questions are already linked to the tables. The test set of the OTT-QA dataset does not have links between the tables with texts. This is out of scope of this current work, but in the future, we plan to explore more in this area.
Acknowledgments
We would like to express our gratitude to Dzmitry Kasinets, a PhD student in the Department of Computer and Information Science and Engineering from University of Florida, for his assistance with the human evaluation of the LLM-generated table-text summaries and independent sub-questions.
Appendix
Appendix A Implementation Details
For all our experiments we have made use of Nvidia Geforce GTX 1660 Ti. We have set the temperature to 0.5 for all experiments that uses zero-shot or few-shot learning. The table-text summary was generated using zero-shot learning with LLaMA 3-70B and for the question decomposition step, we used two-shot learning with LLAMA3B-70B. This was consistent for all our experiments. In the last two stages of our TTQA-RS framework, i.e. for steps involving table-text QA of independent sub-questions and for table-text QA of original questions, we have used few-shot learning with CoT, and we have experimented with different language models such as LLaMA 2-7b, LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, and LLaMA3-70b.
Appendix B Table-text summarization prompt
Figure 6 shows an example of a table-text summarization prompt. The LLM output shows the generated summary.
Figure 6: Prompt for zero-shot table-text summarization
Appendix C Impact of number of shots
In this section, we have performed an ablation study by increasing the number of shots while evaluating our model on the test set of the HybridQA dataset. This is shown in Figure 7. We have evaluated the impact of increasing k in k-shot learning on the baseline standard prompting model, baseline CoT model, and and the TTQA-RS model using LLaMA3 -70B. For standard prompting and CoT, we observe that with an increase in k from 0 to 3, there is an increase in the exact match score. After 3 shots, increasing the number of shots does not improve the performance. For the TTQA-RS model, there is an improvement in EM score from 0-shot to 2-shot, after which increasing the k value does not improve the exact match score of the model.
Figure 7: k-shot ablation study over Hybrid-QA test set

### v2

@@
 Although several works have attempted to solve the table-text QA task, most involve training the models and requiring labeled data.
-In this paper, we have proposed a model - “TTQA-RS: A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization” [^1] Link to code in Github: https://github.com/jayetri/TTQA-RS.
-Our model uses augmented knowledge including table-text summary with decomposed sub-question with answer for a reasoning-based table-text QA.
-Using open-source language models our model outperformed all existing prompting methods for table-text QA tasks on existing table-text QA datasets like HybridQA and OTT-QA’s development set.
-Our results are comparable with the training-based state-of-the-art models, demonstrating the potential of prompt-based approaches using open-source LLMs.
-Additionally, by using GPT-4 with LLaMA3-70B, our model achieved state-of-the-art performance for prompting-based methods on multi-hop table-text QA.
+In this paper, we have proposed a Retrieval Augmented Generation (RAG) based model - “TTQA-RS: A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization” [^1] Link to code in Github: https://github.com/jayetri/TTQA-RS.
+Our model uses an enhanced retriever for table-text information retrieval and uses augmented knowledge, including table-text summary with decomposed sub-questions with answers for a reasoning-based table-text QA.
+Using open-source language models, our model outperformed all existing prompting methods for table-text QA tasks on existing table-text QA datasets, such as HybridQA and OTT-QA’s development set.
+Our experiments demonstrate the potential of prompt-based approaches using open-source LLMs.
+Additionally, by using LLaMA3-70B, our model achieved state-of-the-art performance for prompting-based methods on multi-hop table-text QA.
 TTQA-RS- A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization
@@
 Over the years, several works have attempted to solve this task.
-But the majority of these works have used supervised-training, requiring a large amount of labeled data (Chen et al., 2020b; Sun et al., 2021; Wang et al., 2022; Eisenschlos et al., 2021; Feng et al., 2022; Kumar et al., 2023; Chen et al., 2020a; Li et al., 2021).
+But the majority of these works have used supervised-training for the reader, requiring a large amount of labeled data (Chen et al., 2020b; Sun et al., 2021; Wang et al., 2022b; Eisenschlos et al., 2021; Feng et al., 2022; Kumar et al., 2023; Chen et al., 2020a; Li et al., 2021).
 In this paper, we have proposed a prompting-based approach while using open-source large language models (LLMs) for multi-hop table-text QA.
@@
 models with less than 100B parameters using CoT prompting tend to hallucinate and produce incorrect results, urging research communities to use bigger LLMs which are expensive and also not open-source.
-In this paper, we introduced a framework - TTQA-RS, a reasoning-based prompting approach for table-text QA that despite CoT’s shortcomings on small-parameter models, we were able to reduce the hallucinations on open-source small models (i.e.
-we obtained a 6% increase in exact match score compared to the baseline CoT model for the HybridQA’s test set).
-Furthermore, our proposed model was able to beat the state-of-the-art model - S3HQA’s CoT prompting with GPT 3.5 results (Lei et al., 2023) on HybridQA dataset.
+In this paper, we introduced a framework - TTQA-RS, a prompting approach for table-text QA that despite CoT’s shortcomings on small-parameter models, we were able to reduce the hallucinations on open-source small models.
+Furthermore, our proposed model was able to beat the state-of-the-art model - S3HQA’s CoT prompting using different LLaMA models (Lei et al., 2023) on HybridQA dataset.
 By beating their model’s performance, we have shown the potential for smaller LLMs in multi-hop table-text QA.
@@
 Hence, the OTT-QA’s test set is more challenging.
-Existing models including ours use a retriever-reader framework for table-text QA.
-In this paper, we narrowed our focus to the reader of the table-text QA task.
-Our goal is to develop a prompting strategy for the table-text QA reader that can work even with smaller LLMs.
 The task of linking tables and text passages for open-domain QA is out of scope of this paper.
-The development set of OTT-QA, similar to the HybridQA dataset, already has hyperlinks in the table cells linking to the wiki text.
-In the future, we plan to extend our approach to linking the table and text for cases when hyperlinks are absent in the table cells.
 The TTQA-RS model breaks down the table-text QA problem into multiple steps.
@@
 It also generates the summary of the table and text, which is in turn used for the table-text QA of the original questions.
-Breaking down the complex multi-hop QA problem into simple, smaller steps can help boost the model’s overall performance.
-Furthermore, LLMs struggle with multi-level reasoning in a single step.
-So, breaking down the multi-hop QA problem along with providing an augmented information including the table-text summary can improve the performance of multi-hop table-text QA tasks using small open-source LLMs.
-In Figure 1, we show an example of a question from multi-hop QA that uses standard prompting, CoT, and the TTQA-RS approach for multi-hop QA.
+Breaking down the complex multi-hop QA problem into simple, smaller steps along with providing augmented information, including the table-text summary, can improve the performance of multi-hop table-text QA tasks using small open-source LLMs.
+We have also proposed an enhanced retriever for our framework that further improves the overall performance of our table-text QA model.
+In Figure 1, we show an example of a question that uses standard prompting, CoT, and the TTQA-RS approach for multi-hop QA.
 2 Related Works
@@
 MFORT-QA uses the Chain-of-thought (CoT) method to break down complex questions into smaller sub-questions, and uses Retrieval Augmented Generation to extract more context.
-Similar to the MFORT-QA model, we also break down complex questions into smaller sub-questions.
+Similar to the MFORT-QA model, we also use a RAG-based framework and break down complex questions into smaller sub-questions.
 With the complexity of the multi-hop QA task broken down into smaller questions, LLMs are in turn working on a smaller problem and perform better as single-step reasoners.
-Our model - TTQA-RS, additionally generates a summary using the retrieved table rows and passages.
+Our model - TTQA-RS, additionally uses an enhanced retriever for table-text information retrieval and further generates a summary using the retrieved table rows and passages.
 Then, for table-text question answering (QA), it uses the generated summary, the predicted entity type of the answer, and the generated sub-questions along with the answer.
-Figure 2: An overview of TTQA-RS framework. The dashed lines represent the reader for the table-text QA model.
 3 Our Model
 3.1 System Overview
-The TTQA-RS model uses a retriever-reader model.
+Figure 2: An overview of TTQA-RS framework. The brown dashed lines represent the retriever and the black dotted lines represent the reader for the table-text QA model.
+The TTQA-RS model uses a RAG-based retriever-reader framework.
 Our reader breaks down the table-text QA problem into five steps - (1) Summary generation using retrieved tables rows and passages, (2) Question decomposition, (3) Entity type prediction of the expected answer, (4) Table-text QA of independent sub-question, and (5) Table-text QA of the original question.
 Figure 2 shows an overview of the TTRS-QA framework.
-The following subsections describe the TTQA-RS framework’s retriever and reader in detail.
 3.2 Retriever
-The function of the retriever is to extract relevant rows and passages from the text linked to the table cells using hyperlinks.
-For the HybridQA dataset, we have used S3HQA model’s (Lei et al., 2023) table retriever to extract the relevant row(s) from the table, and HYBRIDER’s (Chen et al., 2020b) text-retriever to extract the relevant information from the linked passages.
-S3HQA’s row retriever uses refinement training to train the retriever model.
-The tables contain hyperlinks to Wikipedia text.
-So, the passages linked to the retrieved rows are collected to form a pool.
-The passage retriever contains an ensemble retriever of TF-IDF retriever with longest-substring retriever and selects passages with cosine distance less than a certain threshold.
-For experiments on OTT-QA’s development set, we don’t use any table retriever, i.e.
-we only use HYBRIDER’s text retriever.
-The text linked to the table rows is extracted and then filtered using HYBRIDER’s text retriever.
+Figure 3: Example of rows to sentences for a table
+We have used an enhanced retriever that first converts rows in the tables to sentences using a pre-trained BART-Large model.
+The questions and sentences are encoded and then fed to a fine-tuned Dense Passage Retriever (DPR) (we can call this the table retriever).
+The table retriever retrievers the top 3 most relevant rows.
+It is important to first convert the rows to sentences using BART since the retriever is not always smart enough to interpret the rows for retrieval.
+This step helps the retriever to make more intelligent decisions.
+Figure 3 shows an example where the rows of the table are converted to sentences.
+In this example, when the table is directly fed to the retriever, it fails to understand that rank 2 implies that the player came second in the National Football League.
+When the rows are converted to sentences before feeding to the DPR, it can better understand the table information and, in turn, retrieve more accurate information.
+For text retrieval, we extract the passages from the text linked to the top 3 rows using hyperlinks.
+The question embedding and row embedding refer to the final dense vectors generated from the encoding process of the question and the rows, respectively.
+We calculate the combined embeddings using the following:
+[…]
+[…]
+Figure 4: Example of our approach using TTQA-RS model
+In the Equation 3.2, $\alpha$ is a parameter with values between 0 and 1.
+We estimate the similarity between the combined embeddings and the text embeddings to retrieve the top 6 candidates using a pre-trained DPR model (i.e.
+the text retriever).
+These top 6 candidates, along with the question keywords generated using a pre-trained T5 model, are fed into a re-ranker, (a pre-trained ’deepset/roberta-base-squad2’ model).
+The re-ranker identifies the top 3 candidates, which are subsequently forwarded to the reader.
 3.3 Reader
@@
 This is the first step of the reader model.
-The retrieved table rows are flattened with a delimiter separating the rows and columns of the table.
 The retrieved rows and passages are used to generate summaries of the table and text.
@@
 In Appendix B we have shown an example of a table-text summarization prompt.
-Figure 3: Example of our approach using TTQA-RS model
 3.3.2 Question decomposition
@@
 We use in-context learning with LLaMA3-70B model to generate the independent sub-questions for the given complex queries.
+Appendix C shows further examples.
 3.3.3 Entity type prediction of the expected answer
@@
 The questions in the HybridQA dataset are based on Wikipedia tables and corpora that are linked to the Wikipedia tables through hyperlinks.
-OTT-QA (Chen et al., 2020a) is an open-domain multi-hop table-text QA dataset.
-For our experiments, we only use the development set of the OTT-QA dataset which contains the hyperlinks linking the table and the text (unlike OTTQA’s test set).
+OTT-QA (Chen et al., 2020a) is an open-domain multi-hop table-text QA dataset. For our experiments, we only use the development set of the OTT-QA dataset.
 4.2 Implementation details
@@
 4.3 Baseline Models
-Standard prompting - For the baseline standard prompting model, we used the same retriever as in TTQA-RS model, (i.e.
-HYBRIDER’s (Chen et al., 2020b) passage retriever with S3HQA’s (Lei et al., 2023) table retriever for the HybridQA dataset).
-For experiments on OTT-QA’s dev set, we don’t use any table-retriever, i.e.
-we only use the HYBRIDER’s passage retriever.
+Standard prompting - For the baseline standard prompting model, we used the same retriever as in TTQA-RS model.
 For the reader, we performed in-context learning with standard prompting (Brown et al., 2020) for the QA task.
-Chain of Thought Prompting (CoT) - Similar to the standard prompting baseline model, the CoT baseline model uses the same retriever as the TTQA-RS model.
-The reader uses in-context learning with CoT prompting Wei et al.
+Chain of Thought Prompting (CoT) - The CoT baseline model also uses the same retriever as the TTQA-RS model. The reader uses in-context learning with CoT prompting Wei et al. (2022).
+Least-to-Most Prompting - Least-to-Most Prompting (LtM) Zhou et al.
+(2022) extends CoT prompting by first diving a problem into subproblems and then solving each one.
+For the baseline LtM model, we use TTQA-RS’s retriever with LtM prompting Zhou et al.
 (2022).
+CoT with Self Consistency (sc) - Wang et al. (2022a) proposed self-consistency as a replacement for the naive greedy decoding utilized in CoT prompting.
 5 Results and Discussion
 5.1 Main Results
+Table 1: Retrieval Accuracy (%)
+Metrics HybridQA OTT-QA
+Dev Test Dev
+HIT@1 73.28 71.49 72.53
+HIT@3 82.01 80.65 81.83
 Table 2: Performance of our model-TTQA-RS and other related works on the HybridQA dataset
@@
 Train DocHopper (Sun et al., 2021) 47.7 / 55.0 46.3 / 53.3
-Train MuGER2 (Wang et al., 2022) 571.1 / 67.3 56.3 / 66.2
+Train MuGER2 (Wang et al., 2022b) 57.11 / 67.3 56.3 / 66.2
 Train POINTR + MATE (Eisenschlos et al., 2021) 63.4 / 71.0 62.8 / 70.2
@@
 Train S3HQA (supervised learning) (Lei et al., 2023) 68.4 / 75.3 67.9 / 75.5
-2-shot S3HQA GPT 3.5 direct (Lei et al., 2023) 57.1 / 68.8 -
-2-shot S3HQA GPT 3.5 CoT (Lei et al., 2023) 60.3 / 72.1 -
-2-shot Baseline Standard prompting LLaMA 3-70B 48.97 / 60.18 52.88 / 61.42
-2-shot Baseline CoT LLaMA 3-70B 54.22 / 64.98 55.71 / 62.24
-2-shot TTQA-RS LLaMA 3 - 70B 63.12 / 73.61 61.97 / 67.56
-2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 65.49 / 76.43 63.69 / 71.83
+2-shot S3HQA LLaMA 3 - 70B CoT (Lei et al., 2023) 52.23 / 58.49 53.75 / 62.94
+2-shot Baseline Standard prompting LLaMA 3-70B 48.73 / 60.02 50.76 / 59.23
+2-shot Baseline CoT LLaMA 3-70B 54.53 / 62.49 55.89 / 62.98
+2-shot Baseline CoT LLaMA 3-70B (with sc) 54.63 / 64.07 55.96 / 63.07
+2-shot Baseline Least-to-Most prompting LLaMA 3-70B 57.84 / 65.42 57.83 / 67.39
+2-shot TTQA-RS LLaMA 3 - 70B 63.97 / 74.20 62.34 / 70.58
+2-shot TTQA-RS LLaMA 3 - 70B (with sc) 63.98 / 74.32 62.92 / 70.83
 Human - 88.2 / 93.5
-(a) EM score
-(b) F1 score
-Figure 4: Performance of HybridQA test set on different LLaMA models
-(a) EM score
-(b) F1 score
-Figure 5: Performance of OTT-QA dev set on different LLaMA models
+Figure 5: EM-score of HybridQA test set on different LLaMA models
 Table 3: Performance of our model - TTQA-RS and other related works on OTT-QA development set.
@@
 Train MITQA (Kumar et al., 2023) 40.0 45.1
-2-shot Baseline Standard prompting LLaMA3-70B 53.28 59.65
-2-shot Baseline CoT LLaMA3-70B 55.74 63.50
-2-shot TTQA-RS LLaMA 3 - 70B 63.15 70.84
-2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 67.27 79.55
+2-shot Baseline Standard prompting LLaMA3-70B 51.24 58.73
+2-shot Baseline CoT LLaMA3-70B 55.78 64.30
+2-shot Baseline CoT LLaMA3-70B (with sc) 55.79 65.54
+2-shot Baseline Least-to-Most prompting LLaMA3-70B 58.68 67.75
+2-shot TTQA-RS LLaMA 3 - 70B 63.17 73.52
+2-shot TTQA-RS LLaMA 3 - 70B (with sc) 63.31 73.54
 Table 4: Ablation studies of TTQA-RS on HybridQA and OTT-QA dataset using LLaMA 3-70B
@@
 Dev Test Dev
-Baseline Standard prompting 48.97 / 60.18 52.88 / 61.24 53.28 / 59.65
-Baseline CoT 54.22 / 64.98 55.71 / 62.24 55.74/63.50
-CoT + entity-type prediction of expected answer 59.68 / 67.47 57.75 / 64.65 57.53 / 66.24
-Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer with COT (no summarization) 61.14 / 70.45 58.72 / 66.40 61.23 / 69.14
-Our model - Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer + summarization with CoT 63.12 / 73.61 61.97 / 67.56 63.15 / 70.84
-Table 4: Human evaluation of generated summaries for a sampled test set of HybridQA
-Human Evaluation Metrics Performance
-Correctness 0.94
-Inclusivity 0.98
-Completeness 0.71
+Baseline Standard prompting 48.73/60.02 50.76/59.23 51.24/58.73
+Baseline CoT 54.53/62.49 55.89/62.98 55.78/64.30
+CoT + entity-type prediction of expected answer 55.34/63.95 56.52/63.47 57.60/67.84
+Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer with COT (no summarization) 60.85/70.68 58.32/66.82 61.06/70.48
+Our model - Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer + summarization with CoT 63.97/74.20 62.34/70.58 63.17/73.52
 In this section, we discuss all our major findings.
+Table 1 shows the hit score of our model on the HybridQA and OTT-QA dataset.
 Table 2 displays the performance of our model with other existing models on the HybridQA dataset.
-We used exact match (EM) and F1 score to evaluate the performance the table-text QA models.
 From Table 2, we can observe that most existing models train their models for table-text QA.
 S3HQA (Lei et al., 2023) is the only model among the existing works that uses in-context learning for the HybridQA dataset.
-Please note that "S3HQA GPT-3.5 direct" refers to S3HQA model with standard prompting using GPT-3.5.
-Our TTQA-RS model with LLaMA 3-70B on the HybridQA’s development set was able to beat S3HQA’s CoT with GPT 3.5 by 3% exact match.
-Our 2-shot model with LLaMA-4 also beats the baseline standard and CoT prompting models by a huge margin (i.e.
-by 9% exact match when compared with standard prompting and by 6% exact match when compared with CoT in the test set).
-Furthermore, in Figure 4 and in Figure 5, we have shown the performance of our TTQA-RS model on different parameter models of LLaMA 2 and LLaMA 3 models on HybridQA test set and the OTT-QA development set respectively.
-For all the different parameter models of LLaMA-2 and LLaMA-3, our framework performed better than the baseline prompting models (i.e.
-standard prompting and CoT prompting).
-Our experiments show that our breakdown prompting approach with summarization and reasoning can improve the performance of all open-source models for table-text QA tasks.
-To show that our TTQA-RS approach can improve table-text QA on also GPT model, we have experimented with GPT-4 in the last stage of our model, i.e.
-table-text QA on the original question.
-For the remaining stages of the reader, we have used LLaMA 3- 70b.
-By adding GPT-4 in the last step, we were able to show the best performance with an exact match of 65.49 and F1 score of 76.43 in the development set, and an exact match of 63.69 and F1 score of 71.83 on the test set.
-With 2-shot learning using TTQA-RS LLaMA 3 70B + GPT-4 model, we were able to reach a model performance very close to the best existing training-based model (S3HQA with supervised learning) on the HybridQA dataset.
-For cost limitations, we have limited experiments with GPT-4 to only the last stage of our model.
+Our TTQA-RS model with LLaMA 3-70B on the HybridQA’s development set was able to beat S3HQA’s CoT model by 11.7% exact match (EM) and by 15.71% F1 score.
+Our model outperformed the S3HQA-CoT model on the HybridQA test using LLaMA3-70B by 8.6% and 7.6% exact match (EM) and F1 score respectively.
+Our 2-shot model with LLaMA 3-70B also beats the baseline standard, baseline CoT, baseline COT with sc, and Least-to-Most prompting method by a huge margin.
+This shows the effectiveness of our reader model.
+We can also observe that our baseline CoT model outperformed the S3HQA’s CoT model for the HybridQA dataset.
+This shows the effectiveness of our retriever model.
+Furthermore, in Figure 5, we have shown the performance of our TTQA-RS model on different parameter models of LLaMA 2 and LLaMA 3 models on HybridQA test set.
+For all the different parameter models of LLaMA-2 and LLaMA-3, our framework performed better than the baseline prompting models and also the S3HQA-CoT model.
+Our experiments show that our approach can improve the performance of open-source models for table-text QA tasks.
 Table 3 shows the performance of our model - TTQA-RS on the OTT-QA development set.
 To the best of our knowledge, MFORT-QA (Guan et al., 2024) is the only model that has used in-context learning for the OTT-QA dataset, but since they have not reported their performance on the development set, we therefore compare our model’s performance with other existing works that trained the models.
-Our TTQA-RS model with LLaMA 3-70B + GPT-4 model achieved the best performance (exact match of 67.27 and F1 score of 79.55) on the development set and has achieved new state-of-the-art performance of the OTT-QA’s development set.
-With the evaluation of our model - TTQA-RS on the HybridQA and OTT-QA development set, we have shown the potential of prompting approaches with small language models (like LLaMA) and also using GPT-4.
+Our TTQA-RS model with LLaMA 3-70B with self-consistency has achieved new state-of-the-art performance on the OTT-QA’s development set.
 5.2 Analysis and Ablation Studies
-This section describes all the analysis and ablation studies performed on our model.
+5.2.1 Effect of converting rows to sentences
+We assessed the retriever model’s effectiveness on the HybridQA test set under two conditions: first, when the table rows are converted into sentences prior to retrieval, and second, when the table rows are used in their original form without conversion.
+We obtained HIT@1 of 71.49 when rows were converted to sentences and HIT@1 of 60.75 when the tables were directly fed to the table retriever.
+This shows that converting the rows to sentences before feeding them to the retriever improves the performance of our table-text QA retriever model.
+5.2.2 Effect of reader
 Table 4 shows the ablation studies of our model using HybridQA dataset and OTT-QA’s development set.
 We can observe that baseline CoT model outperforms the baseline standard prompting model.
-This shows the importance of reasoning in the multi-hop table-text QA task.
-Then, we test the model by adding the entity type prediction of the expected answer in the CoT prompt.
-We notice a significant increase in the performance of the model for all the datasets.
-In the 4th row of Table 3 we have added all the components of our final model except the table-text summary and we can see a further increase in performance in the HybridQA and OTT-QA datasets.
-Finally, we show the performance of our model TTQA-RS (i.e.
-last row) by including the generated table-text summary, and we can observe that our model performs the best with all the steps included (including summarization).
-Adding the table-text summary in the QA input prompt, helps the LLM model to recognize relevant information related to the table or text passages that might have otherwise gone unnoticed.
-This shows the importance of every component of our model and the need for our break-down prompting approach for table-text QA.
-In Appendix  C, we evaluated the impact of the number of shots on the model’s performance.
+We tested the model by adding each component of our reader one by one, and we notice a significant increase in the performance of the model for all the datasets.
+This shows the importance of every component of our model.
+In Appendix  D, we evaluated the impact of the number of shots on the model’s performance.
 5.3 Human Evaluation Results
-We have manually evaluated the first 100 samples of the table-text summaries generated by LLaMA3-70B, and also the independent sub-questions generated using LLM prompting.
+We manually evaluated the first 80 samples of the table-text summaries generated by LLaMA3-70B, and 100 samples of the independent sub-questions generated using LLM prompting were also evaluated.
 We obtained an accuracy of 91% for question decomposition.
-Table  4 tabulated the human evaluation results of the generated summaries for the sampled HybridQA test set.
 For evaluating the generated summaries using retrieved table rows and passages, we have used three evaluation metrics - correctness, inclusivity, and completeness.
@@
 Completeness was used to check if the generated summaries had complete sentences.
-We have included all our human evaluation results in the Supplementary section.
+We obtained a correctness score, inclusivity score, and completeness score of 0.83, 0.87, and 0.98, respectively, for the 80 HybridQA test set samples.
 6 Conclusion
-This paper proposes a prompting strategy of multi-hop table-text QA by generating table-text summaries and answers of sub-questions.
+This paper proposes a RAG-based framework with an enhanced retriever by converting rows to sentences before feeding the tables to a retriever.
+We have further proposed a prompting strategy of multi-hop table-text QA by generating table-text summaries and answers of sub-questions.
 We show that including summaries of retrieved table rows and passages in the prompt with our breakdown approach can substantially increase the performance of CoT prompting in table-text QA.
-The proposed method achieves new state-of-the-art performance among the prompting approaches for multi-hop table-text QA tasks using both open-source (i.e.
-LLaMA3-70B) and GPT-4 models.
-Our experiments specifically focussed on improving prompting strategies in the table-text QA readers.
-In the future, we plan to extend our work on the table-text QA retrievers which can further improve the QA performance.
+The proposed method achieves new state-of-the-art performance among the prompting approaches for multi-hop table-text QA tasks using open-source (i.e.
+LLaMA3-70B) models.
 Limitations
@@
 Errors made in the initial steps can result in wrong answers.
-Furthermore, in the experiment of using GPT-4 on our model, we have limited its usage only to the last step of the reader model as GPT-4 is expensive.
-Using GPT-4 in the remaining steps of the reader could have further improved the performance of our model.
-Secondly, the performance of our prompting-based approach, even though is on par with the fine-tuned state-of-the-art models (or has outperformed the training-based state-of-the-model for OTT-QA development set), it’s performance is still not close to the human performance.
-Also, we are using an existing retriever and the focus of this paper has only been to improve the reader’s performance for multi-hop table-text QA.
-There is still potential to improve the overall performance of the model by using a better table and text retriever for this problem.
+Secondly, the performance of our prompting-based approach, even though is on par with the fine-tuned state-of-the-art models (or has outperformed the training-based state-of-the-model for OTT-QA development set), it’s performance is still not close to the supervised models (in case of HybridQA dataset) and also not close to human performance.
 Also, currently, we have only experimented with multi-hop table-text datasets in which the questions are already linked to the tables.
@@
 This is out of scope of this current work, but in the future, we plan to explore more in this area.
-Acknowledgments
-We would like to express our gratitude to Dzmitry Kasinets, a PhD student in the Department of Computer and Information Science and Engineering from University of Florida, for his assistance with the human evaluation of the LLM-generated table-text summaries and independent sub-questions.
 Appendix
 Appendix A Implementation Details
-For all our experiments we have made use of Nvidia Geforce GTX 1660 Ti.
-We have set the temperature to 0.5 for all experiments that uses zero-shot or few-shot learning.
+For all our experiments, we used Nvidia Geforce GTX 1660 Ti.
+For the retriever framework, we used a pre-trained BART-Large model for converting rows to sentences, with a maximum length set to 50, and the number of returned sequences set to 1.
+For the table retriever, the DPR is fine-tuned with batch size of 32 and trained for 8 epochs, with a learning rate of 2e-5.
+We trained the model for 5000 questions which took 6 hours to train the model.
+For the re-ranker model, we have used the pre-trained ’deepset/roberta-base-squad2’ model with the maximum length set to 512, and number of returned sequences set to 1.
+We have set the $\alpha$ to 0.2 for our experiment.
+For the reader, we have set the temperature to 0.5 for all experiments that use zero-shot or few-shot learning.
 The table-text summary was generated using zero-shot learning with LLaMA 3-70B and for the question decomposition step, we used two-shot learning with LLAMA3B-70B.
@@
 Figure 6: Prompt for zero-shot table-text summarization
-Appendix C Impact of number of shots
+Appendix C Independent sub-questions
+Table  5 shows examples of original complex questions and their corresponding independent sub-questions.
+The sub-answer obtained using the independent sub-question aids in answering the original complex question.
+Table 5: Examples of complex questions and independent sub-questions
+Original Complex Question Independent sub-question
+What was the release date of the game which Andrew Voss provided commentary on ? What game has Andrew Voss provided commentary on ?
+What is the population of the hometown of the 2012 Gatorade Player of the Year ? What is hometown of the 2012 Gatorade Player of the Year ?
+Which battle took place in largest country in sub-Saharan Africa ? What is the largest country in sub-Saharan Africa ?
+What did the Japanese rename the road that Hop Yat Church is located along ? Where is Hop Yat Church located?
+What is the genre of the game that began development as Street Fighter II , but had its name and genre changed ? Which game began development as Street Fighter II, but had its name and genre changed?
+What was the highest legal job of the chancellor of Wilfrid Laurier University after May 1990 ? Who was the chancellor of Wilfrid Laurier University after May 1990?
+What is the availability status of the Buffalo station owned by the E. W. Scripps Company ? Which station in Buffalo is owned by the E. W. Scripps Company?
+Which war film was directed by the director of Hibiscus Town ? Who directed Hibiscus Town?
+How many plays has the person known for The Black Coat written ? Who is the person known for The Black Coat?
+Appendix D Impact of number of shots
 In this section, we have performed an ablation study by increasing the number of shots while evaluating our model on the test set of the HybridQA dataset.